OpenStudy (anonymous):

How many solutions does this equation have ? 2(2x+5)=4(x+3)

OpenStudy (fibonaccichick666):

WHat do you think? What have you tried?

OpenStudy (luckycoins888):

First you need to distribute any numbers outside of the parenthesis through the parenthesis (that was a horrible explanation I'm sorry...) 2(2x+5)=4(x+3) 4x+10=4x+3 then you try to get x on one side of the equation so you can solve for it. 4x-4x=0 10≠3 , therefore there are no solutions to the equation

OpenStudy (luckycoins888):

@Es13 Hope I helped xD